The length of a rectangular room is 1and 2/3 ft more than its width and its perimeter is 46ft.
Find the length and width of the room.

Answer by mananth(16946)

let width be x ft
length will be x+5/3 ft.
..
Perimeter = 2L +2W
=2x+2(x+5/3)
=2x+2x+10/3
=4x+(10/3) =46
=12x+10 /3 = 46
12x +10 =138
12x= 138-10
12x=128
x= 10.66 feet width
length = 10.66+1.75 = 12.41 feet